OpenStudy (anonymous):

Find the 31st term of the following sequence. 9, 15, 21, ... 186 189 195 is there an easy way to do this

OpenStudy (anonymous):

Find the nth term for this sequence using Tn = (a2 - a1)n + 2a1 - a2 where a1 is the first term and a2 is the second term, then substitute n = 31

OpenStudy (anonymous):

3+6n

OpenStudy (jhannybean):

the common difference is found by subtracting the second number from the first: i.e 21-15, 15-9, etc...What do you get as the common difference?

OpenStudy (mathstudent55):

First, can you tell if this is an arithmetic or geometric sequence?

OpenStudy (anonymous):

tn=a+(n-1)d here d=15-9=6 a=9 n=31 plug the values and get the solution

OpenStudy (anonymous):

to complicated nvm byby

OpenStudy (anonymous):

nvm im back explain again plz

OpenStudy (anonymous):

t31=9+(31-1)*6 solve

OpenStudy (anonymous):

everyone give me a medal thers like a 10 people in here lol

OpenStudy (jhannybean):

As what @mathstudent55 was saying, when you look at a sequence of numbers you want to see whether the sequence is geometric \(\large a_n=a_1r^{(n-1)}\), or arithmetic \(\large a_n = a_1 +(n-1)d\), by checking to see whether it has a common difference \(\large a_{n-1} - a_n\) or if has a common ratio, \(\large \frac{a_{n-1}}{a_n}\). Once found, you can easily deduce what type of formula to apply your understanding of the sequence to.

OpenStudy (anonymous):

Arithmetic sequence can be of higher tier so you might want to be specific with your formulas. an = a1 + (n - 1)d is for linear sequence

OpenStudy (jhannybean):

Elaborate?

OpenStudy (anonymous):

You can tell if a sequence is linear by comparing a2 - a1 with a3 - a2 and both of them are not 0

OpenStudy (jhannybean):

Ok, yes you are correct.

OpenStudy (anonymous):

There's quadratic sequence, cubic etc

OpenStudy (jhannybean):

I don't believe this person is that far into understanding these though.. Maybe im wrong! :P

OpenStudy (anonymous):

so its 189 ??????????????????????????????? thts what i got

OpenStudy (anonymous):

It's better to clarify early than late

OpenStudy (anonymous):

Yes 189

OpenStudy (jhannybean):

You got it @dirtydan667

OpenStudy (anonymous):

yayayyayay

OpenStudy (jhannybean):

You took it away? :(